how do the y-coordinates change from (7,10) to (7,-2) in this translation?
Answer by Edwin McCravy(20065)



The upper black dot is at (7,10)
The lower black dot is at (7,-2)
The upper black dot at (7,10) must fall twelve units to be where the
lower black dot is. A fall of twelve units is indicated by -12.
Here's how to get that
end point's y-coordinate, -2, minus beginning point's y-coordiate, +10,
(-2)-(10) = -2-10 = -12
The change in the y coordinate is -12.
Edwin